KAMI/KTeV DAY Shift Summary -- 1/19/2000 (Wed.)


Subject: KAMI/KTeV DAY Shift Summary -- 1/19/2000 (Wed.)
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:32:01 CST


              Shift Summary for day shift, Wed. 1/19/2000

On Shift: MT, TA and RK + ...

Hours of beam available at all intensities: 8.5
Hours of beam available at nominal intensity: 8.5
Hours of KAMI test data: 7.5
Hours of special runs (laser scans, muon runs...): 0.0

 
Accounting of lost time & intensity:

  starting/stopping runs
 
========================================================================
Runs ended this shift:

       15612 to 15619 -- test KAMI_HADRON
       Tapes: QKE210-218, QKB170 for runs 15612-15619

========================================================================
Tapes brought to FCC this shift:

       None.

========================================================================
Accesses during shift:

       None

========================================================================
Overview of shift:

       o KAMI test running in HADRON mode.
         Measurement of the neutron/kaon ratio
         (using various absorber thickness) with the UCAL.
       o Measure the kaon flux using KL->3pi0

========================================================================
Description of Shift: from ~da_ktev/doc/runkami.log

                Summary of n,k Flux runs @ 800 GeV(Jan 19)
               ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

       C07: analog sum threshold @ 821, 60 mV ~ 1 MIP
       C09: >=3 hits above 4413 thresh, NMIP= (thr-20mV)/40

                                               C07 C09 RAW C07
                  3pi0-PS: UCAL absorber thr thr Y3p0 rate
   RUN <SEM> B06 C11,13 PS config mV mV in C11 KHz
   ------------------------------------------------------------------------
   15612 2E12 5 10? 60,000 20" Be 50 60 55.3 K 890 KHz
   15614 5E11 1 1,1 2,000 20" Be 50 100 24.1 K 205
   15615 5E11 1 1,1 2,000 20" Be 50 100 47.1 K 230

   15616 5E11 1 1,1 2,000 10" Be 50 100 77.6 K 390

   15617 20" Be (ignore; wanted shadow too)
   15618 2.5E11 1 1,1 2,000 0" Be 50 100 64.7 K 323
   15619 4.7E11 1 1,1 2,000 0" Be 50 100 13.1 K 710

   ------------------------------------------------------------------------
   The B06 vac PS was 2/5 for runs 15612-4: this leads to slight
   loss of events since a few % are recorded as vac ??

   
   Summary of on-line n/k ratios (using ucal#nkratio):

                       on-line
            Absorber n/k
          ----------------------------
               0" Be 1.51
              10" Be 1.13
              20" Be 1.00 (by defintion)
              38" Be --> will come tonight
              ?? Wax --> will come tonight
          ----------------------------

========================================================================
Follow up Notes on Previous Shift:

 keep collecting AGEL data on floppy disks.

========================================================================
Items to Follow up on Next Shifts:
 
 keep collecting AGEL data on floppy disks.

========================================================================



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:32:05 CST